A 1 kg poodle running right at 3 m/s has an elastic head on collision with a 9 kg basset hound walking left with a velocity of 1 m/s. After the collision the poodle bounces off to the left at 3 m/s. Whats the basset hounds velocity after the collision?
1
Expert's answer
2012-09-25T11:15:20-0400
Let m1 is the poodle's weight and V1 is its velocity, m2 is the
basset
hound's weight and V2 is its velocity before the collision.
m1 * V1
- m2 * V2 = - m1 * V1' + m2 * V2'
V2' = (m1 * V1 - m2 * V2 + m1 * V1') /
m2
V2' = (1 * 3 - 9 * 1 + 1 * 3) / 9 = - 1/3 m/s
Answer: 1/3 m/s (directed
to the left).